# Coreset Strikes Back: Improved Parameterized Approximation Schemes for (Constrained) $k$-Median/Means

### v2

@@ 1.2 Technical Overview
Finally, we note that a recent work of Huang et al. [HLLW25] gives a coreset framework for a number of constrained $(k,z)$-clustering objectives, including capacities and fairness, in several metric spaces. For continuous Euclidean, planar, and minor-free metrics, they design coresets for these objectives whose size is independent of $\log n$. In Section 4.4, we sketch an approach to obtain EPASes in the corresponding settings via essentially brute forcing on such coresets. However, this approach results in substantially worse running times as compared to our approach. Furthermore, although their framework designs coresets for many metric spaces that have bounded scatter dimension; it does not appear to give such a result for all metric spaces of bounded scatter dimension. For example, metrics of bounded highway dimension (which have bounded scatter dimension), do not seem to be captured by their framework. Therefore, in our opinion, their result along with brute force enumeration does not satisfactorily resolve 1.
@@ 2 Preliminaries
We consider metric (clustering) space $M=(P,F,{\sf d})$, where $P$ is a set of $n$ points/clients, and $F$ is a (possibly infinite) set of facilities, and ${\sf d}$ is a metric over $P\cup F$. Furthermore, a class $\mathcal{M}$ of metric spaces is a potentially infinite set of metric spaces. Note that, in case $F$ is infinite, the set $F$ may not be explicitly provided in the input.
For a real $B>1$ and $\delta>0$, we denote by $[B]_{\delta}$, the set containing $(1+\delta)^{j}$, for integers $0\leq j\leq\frac{2\log B}{\log(1+\delta)}$.
@@ 2.1 Scatter Dimension
In this section, we revise useful facts related to algorithmic scatter dimension. For a comprehensive background on scatter dimension, see [ABB+23].
@@
In the Ball intersection problem, we are given a metric space $M=(P,F,{\sf d})$ from a metric class $\mathcal{M}$, a finite set $Q\subsetneq P\times\mathbb{R}_{\geq 0}$ of point-distance constraints, called request set, and a an error parameter $\eta>0$. The goal is to find a center $x\in F$ such that ${\sf d}(p,x)\leq(1+\eta)\alpha$, for all $(p,\alpha)\in Q$, if such center exists, and report failure ()bottom($\bot$)( ⊥ ) otherwise.
@@
[…]
  • $x_{i}=\texttt{BALL-INT}_{\mathcal{M}}(M,\{(p_{1},\alpha),\dots,(p_{i-1},\alpha_{i-1})\},\varepsilon/2)$$\forall 2\leq i\leq\kappa$
@@
[…]
The following lemma is handy for bounding the number of iterations of our algorithm.
Lemma 9 [ABB+23, GI25]).
For any class $\mathcal{M}$ of metric space with algorithmic $\varepsilon$-scatter dimension $\lambda(\varepsilon)$, there exists a BALL-INT algorithm $\mathcal{C}_{\mathcal{M}}$ with the following property. Given $\varepsilon\in(0,1)$, a constant $t\geq 1$, and $a_{i}>0,\tau_{i}\geq 2$ for $i\in[t]$, any $(\mathcal{C}_{\mathcal{M}},\varepsilon)$-scattering contains $O(\sum_{i\in[t]}\lambda(\nicefrac{{\varepsilon}}{{2}})(\log\tau_{i})/\varepsilon)$ many triples whose radii lie in the interval $\cup_{i\in[t]}[a_{i},\tau_{i}a_{i}]$.
@@ 3 Framework
In this paper, we consider $(k,z)$-clustering problems with assignment constraints, called Constrained $(k,z)$-Clustering. The general setup is as follows. Let $z\geq 1$ be a fixed integer. We are given an instance $\mathcal{I}$ of Constrained $(k,z)$-Clustering over a metric space $M=(P,F,{\sf d})$, and an integer $k\geq 1$. In addition, there may be problem-specific constraints (such as capacities, fairness, etc.). The goal is to select a set $X\subseteq F$ of size $k$ and an assignment $f:P\to X$, ensuring that both satisfy the given problem-specific constraints that could potentially depend on $X$. We refer to such $X$ as a feasible set of centers and such an $f$ as a feasible assignment from $P$ to $X$ The cost of pair $(X,f)$ — which we also refer to as a solution—is defined as ${\sf cost}(P,X,f)\coloneqq\sum_{p\in P}{\sf d}(p,f(p))^{z}$. Furthermore, the cost of a feasible set of $k$ centers $X\subseteq F$ is defined as ${\sf cost}(P,X)\coloneqq\min_{f}{\sf cost}(P,X,f)$, where the minimum is taken over all feasible assignments $f$. If no feasible assignment $f:P\to X$ exists, or if $X$ is not feasible, then we define ${\sf cost}(P,X)\coloneqq\infty$. The goal is to find a solution $(X,f)$ that minimizes the cost of the solution. In addition, our framework requires that the Constrained $(k,z)$-Clustering problem satisfies the following properties.
@@
Consider a subset of clients $Y\subseteq P$, and a weight function $w:Y\to\mathbb{N}$. For a $k$-sized subset $C\subseteq F$ of centers, we say that $f:Y\times C\to\mathbb{N}$ is a feasible assignment from $Y$ to $C$ if (i) for each $p\in Y$, $\sum_{c\in C}f(p,c)=w(p)$, and (ii) $f$ satisfies additional problem-specific constraints [^10] Note that this definition of a feasible assignment naturally generalizes the definition of an assignment from an unweighted set $P$ to $X$, given earlier: if $f:P\to X$ is a feasible assignment, then construct an assignment $f^{\prime}:P\times X\to\left\{0,1\right\}$ by letting $f^{\prime}(p,c)=1$ iff $f(p)=c$. . We define the weighted-cost of $(Y,C,f)$ as ${\sf wcost}(Y,X,f)\coloneqq\sum_{p\in Y}\sum_{c\in C}f(p,c)\cdot{\sf d}(p,c)^{z}$. As before, we define the weighted-cost of $X$ as ${\sf wcost}(Y,C)\coloneqq\min_{f}{\sf wcost}(Y,C,f)$, where the min is taken over all feasible assignments $f:Y\times C\to\mathbb{N}$, and if no such feasible assignment exists, then ${\sf wcost}(Y,C,f)\coloneqq\infty$.
For an $0<\varepsilon\leq 1$, we say that $(Y,w)$ is an $\varepsilon$-coreset if, for each $C\subseteq F$ of size $k$, it holds that $|{\sf wcost}(Y,C)-{\sf cost}(P,C)|\leq\varepsilon\cdot{\sf cost}(P,C)$, with high probability.
We require that, for any $10<\varepsilon<$ 1, an $\varepsilon$-coreset for the Constrained $(k,z)$-clustering problem of size $\left(\frac{k\log n}{\varepsilon}\right)^{O(1)}$ is computable in polynomial-time in general metric spaces. We also allow the size of the coreset to depend on some problem-specific parameters (e.g., the number of colors in fair clustering), which will be reflected in the final running time.
@@
There exists an algorithm, ASSIGNMENT which takes as input a weighted point set $(Y,w)$, a candidate set $X\subseteq F$ of $k$ centers, and a parameter $10<\varepsilon<$ 1, and returns a feasible assignment $f:Y\times X\to\mathbb{N}$ such that ${\sf wcost}(Y,X,f)\leq(1+\varepsilon)\cdot{\sf wcost}(Y,X)$, where recall that ${\sf wcost}(Y,X)$ is the minimum-cost assignment from $Y$ to $X$. If no such assignment exists, then the algorithm returns $\bm{\bot}$. We require that the algorithm runs in time FPT in $k,\varepsilon$, and other problem-specific parameters. Note that, when this algorithm is used with the original (unweighted) point set $P$ with unit-weight function $\mathbbm{1}$, it returns a feasible assignment $f:P\times X\to\left\{0,1\right\}$, which can be interpreted as a feasible assignment $f:P\to X$ in a natural way. Thus, it suffices to compute a set of centers $X$, since a near-optimal assignment from either coreset $(Y,w)$ or the original point set $P$ to $X$ can be computed using ASSIGNMENT algorithm at the overhead of FPT time. We denote by $T_{\textsf{A}}(\mathcal{I})$ as the running time of ASSIGNMENT on input instance $\mathcal{I}$ [^11] For simplicity, we hide the dependence of the running time on $\varepsilon$.
@@ 3.2 Algorithm
The pseudo-code of our main algorithm yielding Theorem 10 is described in Algorithm 1. This algorithm is based on the ideas mentioned in Section 1.1, and it uses the recursive algorithm Algorithm 2 as a subroutine. It takes as input, an instance $\mathcal{I}$ of Well Constrained $(k,z)$-Clustering, a ball intersection algorithm BALL-INT, an assignment algorithm ASSIGNMENT, and a guess $\mathcal{G}$ of optimal cost. Note that the guess, $\mathcal{G}$ of optimal cost, can be computed efficiently, up to a multiplicative factor of $(1+\varepsilon)$.
@@
1:Input: Instance $\mathcal{I}$ of Well Constrained $(k,z)$-Clustering in metric space $M=(P,F,{\sf d})$, $\varepsilon\in(0,1/2]$, Ball intersection algorithm BALL-INT, ASSIGNMENT Algorithm and ${\sf OPT}\leq\mathcal{G}\leq(1+\varepsilon)\cdot{\sf OPT}$.
@@
6:     Let =$\mathcal{B}=\{B_{1},\dots,B_{k}\}$, where $B_{i}={\sf ball}(p^{\prime}_{i},\nicefrac{{6r^{\prime}_{i}}}{{\varepsilon}})$
@@
1:Let $X=(x_{1},\dots,x_{k})$ such that $x_{i}=\texttt{BALL-INT}(Q_{i},\varepsilon),i\in[k]$
2:return $\bot$ if no such $X$ was found
3:$f\leftarrow\texttt{ASSIGNMENT}(Y,X,(Q_{1},\dots,Q_{k}),\varepsilon)$, return $\bot$ if it fails
4:If ${\sf cost}(Y,X,f)\leq(1+5\varepsilon)\cdot\mathcal{G}$ then return $X$
@@
[…]

### v3

@@ 1.4.2 Challenges of Extending to Algorithmic Scatter Dimension
We note that a recent work of Huang et al. [HLL+25] gives a coreset framework for a number of constrained $(k,z)$-clustering objectives, including capacities and fairness, in several metric spaces. For continuous Euclidean, planar, and minor-free metrics, they design coresets for these objectives whose size is independent of $\log n$. In Section 4.3, we sketch an approach to obtain EPASes in the corresponding settings via essentially brute forcing on such coresets. However, this approach results in substantially worse running times as compared to our approach (e.g., $2^{\Omega_{\varepsilon}(k^{3}\log k)}\cdot\mathcal{I}^{O(1)}$ vs. $2^{O_{\varepsilon}(k\log k)}\cdot\mathcal{I}^{O(1)}$, see  Section 4.1.1 for Capacitated setting). Furthermore, although their framework designs coresets for many metric spaces that have bounded scatter dimension; it does not appear to give such a result for all metric spaces of bounded scatter dimension. For example, metrics of bounded highway dimension (which have bounded scatter dimension), do not seem to be captured by their framework. Therefore, in our opinion, their result along with brute force enumeration does not satisfactorily resolve 1.
This paper Present
FPT factor $2^{O_{\varepsilon}(k\log k)}$ $2^{O_{\varepsilon}(k^{3}\log k)}$ (follows from [HLL+25])
Table 2: EPAS running time for Uniform capacitated $k$-Median in high dimensional Euclidean spaces.
@@ 2 Preliminaries
We consider a metric (clustering) space $M=(P,F,{\sf d})$, where $P$ is a set of $n$ points/clients, and $F$ is a (possibly infinite) set of facilities, and $\sf d$} is a metric over $P\cup F$. Let $\Delta_{M}$ be the distance aspect ratio of $M$; for brevity we use $\Delta$ instead of $\Delta_{M}$ when $M$ is clear from the context. Furthermore, a class $\mathcal{M}$ of metric spaces is a potentially infinite set of metric spaces. Note that, in case $F$ is infinite, the set $F$ may not be explicitly provided in the input. Additionally, points in $P$ have weights defined by the weight function $w:P\rightarrow\mathbb{R}_{\geq 0}$.
For a real $B>1$ and $\delta>0$, we denote by $B]_{\delta}$, the set containing $(1+\delta)^{j}$, for integers $0\leq j\leq\frac{2\log B}{\log(1+\delta)}$.
@@ 2.1 Scatter Dimension
In this section, we revise useful facts related to scatter dimension that were introduced in [ABB+23]. For a comprehensive background on scatter dimension, see [ABB+23].
Definition 5 ($\varepsilon$-Scatter Dimension).
Given a class $\mathcal{M}$ of finite metric spaces, a metric space $M=(P,F,d)\in\mathcal{M}$, and $\varepsilon\in(0,1)$, an $\varepsilon$-scattering sequence in $M$ is a sequence $(x_{1},p_{1}),\cdots,(x_{\ell},p_{\ell})$ of center-point pairs $(x_{i},p_{i})\in F\times P$ such that $d(x_{i},p_{i})>(1+\varepsilon)$ for all $i\in[\ell]$, and $d(x_{j},p_{i})\leq 1$ for all $1\leq i<j\leq\ell$. The $\varepsilon$-scatter dimension of $M$ is the length of the longest $\varepsilon$-scattering sequence present in $M$. The $\varepsilon$-scatter dimension of $\mathcal{M}$ is the supremum of the $\varepsilon$-scatter dimension over all $M\in\mathcal{M}$.
In this paper, we consider the following equivalent definition of $\varepsilon$-scatter dimension, that is helpful in analyzing the algorithms.
Definition 6 ($\varepsilon$-Scatter Dimension (alternate definition)).
Given a class $\mathcal{M}$ of finite metric spaces, a metric space $M=(P,F,d)\in\mathcal{M}$, and $\varepsilon\in(0,1)$, an $\varepsilon$-scattering sequence in $M$ is a sequence $(x_{1},p_{1},\alpha_{1}),\cdots,(x_{\ell},p_{\ell},\alpha_{\ell})$ of center-point-radius triples $(x_{i},p_{i},\alpha_{i})\in F\times P\times\mathbb{R}_{\geq 0}$ such that $d(x_{i},p_{i})>(1+\varepsilon)\alpha_{i}$ for all $i\in[\ell]$, and $d(x_{j},p_{i})\leq r_{i}$ for all $1\leq i<j\leq\ell$. The $\varepsilon$-scatter dimension of $M$ is $\lambda_{M}(\varepsilon)$ if any $\varepsilon$-scattering sequence contains at most $\lambda_{{M}}(\varepsilon)$ many triples per radius value. The $\varepsilon$-scatter dimension of $\mathcal{M}$ is the supremum of the $\varepsilon$-scatter dimension over all $M\in\mathcal{M}$.
@@
In the Ball intersection problem, we are given a metric space $M=(P,F,{\sf d})$ from a metric class $\mathcal{M}$, a finite set $Q\subsetneq P\times\mathbb{R}_{\geq 0}$ of point-distance constraints, called request set, and a an error parameter $\eta>0$. The goal is to either (i) find a center $x\in F$ such that for all $(p,\alpha)\in Q$ if ${\sf d}(p,x)\leq(1+\eta)\alpha$; or (ii) report failure indicating that there does not exist a center $x^{\star}\in F$ such that for all $(p,\alpha)\in Q$, ${\sf d}(p,x^{\star})\leq\alpha$.
@@
Let $\mathcal{M}$ be a class of metric spaces with ball intersection algorithm $\texttt{BALL-INT}_{\mathcal{M}}$. Let $M\in\mathcal{M}$ and $\varepsilon\in(0,1)$. A $(\texttt{BALL-INT}_{\mathcal{M}},\varepsilon)$-scattering sequence is a sequence $(x_{1},p_{1},\alpha_{1}),\dots,(x_{\kappa},p_{\kappa},\alpha_{\kappa})$, where $\kappa$ is some positive integer, and for $i\in[\kappa],x_{i}\in F$, $p_{i}\in P$ and $\alpha_{i}\in\mathbb{R}_{+}$ such that
  • $x_{i}=\texttt{BALL-INT}_{\mathcal{M}}(M,\{(p_{1},\alpha_{1}),\dots,(p_{i-1},\alpha_{i-1})\},\varepsilon/2)$$\forall 2\leq i\leq\kappa$
@@
The algorithmic $(\varepsilon,\texttt{BALL-INT}_{\mathcal{M}})$-scatter dimension of $\mathcal{M}$ is $\lambda_{\mathcal{M}}(\varepsilon)$ if any $(\texttt{BALL-INT}_{\mathcal{M}},\varepsilon)$-scattering sequence contains at most $\lambda_{\mathcal{M}}(\varepsilon)$ many triples per radius value. The algorithmic $\varepsilon$-scatter dimension of $\mathcal{M}$ is the minimum algorithmic $(\varepsilon,\texttt{BALL-INT}_{\mathcal{M}})$-scatter dimension over any ball intersection algorithm $\texttt{BALL-INT}_{\mathcal{M}}$ for $\mathcal{M}$.
The following lemma from [ABB+23] relates $\varepsilon$-scatter dimension of a class $\mathcal{M}$ of finite metric spaces to its algorithmic $\varepsilon$-scatter dimension. The proof constructs a ball intersection algorithm that simply searches for a feasible center exhaustively. This allows us to bound $(\varepsilon,\texttt{BALL-INT})$-scatter dimension of $\mathcal{M}$ for any ball intersection algorithm BALL-INT admitted by $\mathcal{M}$.
Lemma 9 ([ABB+23]).
Every class $\mathcal{M}$ of finite and explicitly given metric spaces with bounded $\varepsilon$-scatter dimension $\lambda(\varepsilon)$ has algorithmic $\varepsilon$-scatter dimension $\lambda(\varepsilon)$. In fact, for any ball intersection algorithm BALL-INT admitted by $\mathcal{M}$, algorithmic $(\varepsilon,\texttt{BALL-INT})$-scatter dimension of $\mathcal{M}$ is bounded by $\lambda(\varepsilon)$.
In this paper, we design an EPAS framework for $(k,z)$-Clustering in metric spaces of bounded $\varepsilon$-scatter dimension. But the analysis of our framework (and algorithms) is based on its algorithmic $\varepsilon$-scatter dimension, which is also bounded due to the above lemma. Towards this, we will construct a suitable ball intersection algorithm that guarantees the constraints of the problem are satisfied. For this, we use the following lemma that is handy for bounding the number of iterations of our algorithm.
Lemma 10 ([ABB+23, GI25]).
For any class $\mathcal{M}$ of metric spaces with algorithmic $\varepsilon$-scatter dimension $\lambda(\varepsilon)$, there exists a BALL-INT algorithm $\mathcal{C}_{\mathcal{M}}$ with the following property. Given $\varepsilon\in(0,1)$, a constant $t\geq 1$, and $a_{i}>0,\tau_{i}\geq 2$ for $i\in[t]$, any $(\mathcal{C}_{\mathcal{M}},\varepsilon)$-scattering contains $O(\sum_{i\in[t]}\lambda(\nicefrac{{\varepsilon}}{{2}})(\log\tau_{i})/\varepsilon)$ many triples whose radii lie in the interval $\cup_{i\in[t]}[a_{i},\tau_{i}a_{i}]$.
@@ 3 Framework
In this paper, we consider $(k,z)$-Clustering with assignment constraints, called constrained $(k,z)$-Clustering. The general setup is as follows. Let $z\geq 1$ be a fixed integer. We are given an instance $\mathcal{I}$ of constrained $(k,z)$-Clustering over a metric space $M=(P,F,{\sf d})$ along with weight function $w:P\rightarrow\mathbb{R}_{\geq 0}$, and an integer $k\geq 1$. The goal is to select a set $X\subseteq F$ of size $k$ and an assignment $f:P\times X\rightarrow\mathbb{R}_{\geq 0}$ with $\sum_{x\in X}f(p,x)=w(p)$ for every $p\in P$, ensuring that $(X,f)$ is feasible. We refer to such $X$ as a feasible set of centers and such an $f$ as a feasible assignment from $P$ to $X$. The cost of pair $(X,f)$ — which we also refer to as a solution—is defined as ${\sf cost}(P,X,f)\coloneqq\sum_{p\in P}\sum_{x\in X}f(p,x)\cdot{\sf d}(p,x)^{z}$. Furthermore, the cost of a feasible set of $k$ centers $X\subseteq F$ is defined as ${\sf cost}(P,X)\coloneqq\min_{f}{\sf cost}(P,X,f)$, where the minimum is taken over all feasible assignments $f$. If $(f,X)$ is feasible, we refer to $(X,f)$ as a feasible solution. If no feasible assignment $f$ exists, or if $X$ is not feasible, then we define ${\sf cost}(P,X)\coloneqq\infty$. The goal is to find a feasible solution $(X,f)$ with minimum cost. In addition, our framework requires that the Constrained $(k,z)$-Clustering problem satisfies the following properties.
@@
Consider a subset of clients $Y\subseteq P$, and a weight function $w:Y\to\mathbb{N}$. For a $k$-sized subset $C\subseteq F$ of centers, we say that $f:Y\times C\to\mathbb{R}_{\geq 0}$ is a feasible assignment from $Y$ to $C$ if (i) for each $p\in Y$, $\sum_{c\in C}f(p,c)=w(p)$, and (ii) $f$ satisfies additional problem-specific constraints [^13] Note that this definition of a feasible assignment naturally generalizes the definition of an assignment from an unweighted set $P$ to $X$, given earlier: if $f:P\to X$ is a feasible assignment, then construct an assignment $f^{\prime}:P\times X\to\left\{0,1\right\}$ by letting $f^{\prime}(p,c)=1$ iff $f(p)=c$. . We define the weighted-cost of $(Y,C,f)$ as ${\sf wcost}(Y,X,f)\coloneqq\sum_{p\in Y}\sum_{c\in C}f(p,c)\cdot{\sf d}(p,c)^{z}$. As before, we define the weighted-cost of $X$ as ${\sf wcost}(Y,C)\coloneqq\min_{f}{\sf wcost}(Y,C,f)$, where the min is taken over all feasible assignments $f:Y\times C\to\mathbb{N}$, and if no such feasible assignment exists, then ${\sf wcost}(Y,C,f)\coloneqq\infty$.
For an $0<\varepsilon\leq 1$, we say that $(Y,w)$ is an $\varepsilon$-coreset if, for each $C\subseteq F$ of size $k$, it holds that $|{\sf wcost}(Y,C)-{\sf cost}(P,C)|\leq\varepsilon\cdot{\sf cost}(P,C)$.
We require that, for any $0<\varepsilon<1$, an $\varepsilon$-coreset for the Constrained $(k,z)$-clustering problem of size $\left(\frac{k\log n}{\varepsilon}\right)^{O(1)}$ is computable in polynomial time [^14] For randomized algorithms, we want that the algorithm outputs an $\varepsilon$-coreset with high probability. in general metric spaces. We also allow the size of the coreset to depend on some problem-specific parameters (e.g., the number of colors in fair clustering), which will be reflected in the final running time.
@@
[…]
@@
[…]
  • The returned center set $X$ must be a feasible.
  • Compatibility with ASSIGNMENT. Let $X$ be set of $k$ centers obtained from GenBALL-INT satisfying request sets $Q_{1},\dots,Q_{k}$, and let $f$ be a near-optimal assignment obtained from ASSIGNMENT subroutine for $X$. Then, for every feasible solution $(Y,g)$ for the problem, there exists a bijection $b:Y\leftrightarrow X$ such that that $(X,g^{\prime})$ is also a feasible solution, where $g^{\prime}(p,x):=g(p,f(y))$, for $p\in P,x\in X$.
@@ 3.2 Algorithm
The pseudo-code of our main algorithm yielding Theorem 11 is described in Algorithm 1. This algorithm is based on the ideas mentioned in Section 1.3, and it uses the recursive algorithm Algorithm 2 as a subroutine. Algorithm 1 takes as input, an instance $\mathcal{I}$ of Well Constrained $(k,z)$-Clustering, a ball intersection algorithm BALL-INT, an assignment algorithm ASSIGNMENT, and a guess $\mathcal{G}$ of optimal cost. Note that the guess, $\mathcal{G}$ of optimal cost, can be computed efficiently, up to a multiplicative factor of $(1+\varepsilon)$. The main task of the subroutine Algorithm 2 is to compute a feasible solution $(X,f)$ for the set of requests $(Q_{1},\dots,Q_{k})$ it receives, where $X$ is a set of centers and $f$ is a feasible assignment of clients to those centers. If the resulting solution has cost close to the guessed optimum, the algorithm terminates and returns it. Otherwise, it identifies an unhappy pair $(p,i)$ with respect to $(X,f)$, modifies the request set $Q_{i}$ accordingly, and recurses.
@@
1:Input: Instance $\mathcal{I}$ of Well Constrained $(k,z)$-Clustering in metric space $M=(P,F,{\sf d})\in\mathcal{M}$, $\varepsilon\in(0,1/2]$, Ball intersection algorithm BALL-INT, ASSIGNMENT Algorithm and ${\sf OPT}\leq\mathcal{G}\leq(1+\varepsilon)\cdot{\sf OPT}$.
@@
6:  Let $\mathcal{B}=\{B_{1},\dots,B_{k}\}$, where $B_{i}={\sf ball}(p^{\prime}_{i},\nicefrac{{6r^{\prime}_{i}}}{{\varepsilon}})$
@@
Algorithm 2 $\texttt{COMPUTE-SOLUTION}((Q_{1},\dots,Q_{k}),\mathcal{B})$
1:Let $X=(x_{1},\dots,x_{k})\leftarrow\texttt{GenBALL-INT}(Q_{1},\dots,Q_{k})$
2:return $\bot$ if no such $X$ was found
@@
4:If ${\sf wcost}(Y,X,f)\leq(1+5\varepsilon)\cdot\mathcal{G}$ then return $X$
@@
6:  return $\texttt{COMPUTE-SOLUTION}((Q_{1},\dots,Q_{i}\cup\{(p,\frac{{\sf d}(p,x_{i})}{1+\varepsilon})\},\dots,Q_{k}),\mathcal{B})$ if not $\bot$